I need help with this problem. Please qnd thank you!
The units' digit of a two digit number is 4 less than the tens' digit. the number it names is 3 less than 9 times the sum of the digits. what is the number?

Found 2 solutions by mathstutor494, lwsshak3:
Answer by mathstutor494(120)

Question is incorrect. Pls check & correct syntax.

Answer by lwsshak3(11628)

The units' digit of a two digit number is 4 less than the tens' digit. the number it names is 3 less than 9 times the sum of the digits. what is the number?
***
let u=units digit
let t=tens digit
..
u=t-4
10t+u=9(t+u)-3
10t+t-4=9t+9u-3
10t+t-4=9t+9(t-4)-3
10t+t-4=9t+9t-36-3
11t-4=18t-36-3
7t=35
t=5
u=t-4=1
number:51